OpenStudy (anonymous):

Ricky took 30 minutes to wash the car. Derek took 50 minutes. How long would it take them to wash the car altogether?

OpenStudy (anonymous):

1 hour and 20 minutes

OpenStudy (anonymous):

1/30m + 1/50m=150

OpenStudy (australopithecus):

why would it take a longer time for 2 people to wash a car? That doesn't make any sense logically.

OpenStudy (australopithecus):

Although My reasoning is probably wrong. I'm on 2 hours of sleep doing math till I pass out

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

sorry thats not right, you cant just subtract their two rates to find 1 completed

OpenStudy (anonymous):

fonzy is closer, but i think you set it =1 to one, i dont remember if the fractions are setup correctly or not though

OpenStudy (anonymous):

See how problems like this one on http://www.saab.org/mathdrills/workcompletion.cgi

OpenStudy (anonymous):

Let x be the time needed for both to wash the car then \[ \frac{1}{30}+\frac{1}{50}=\frac{1}{x} \] If the first can finish it in 30 minutes so in one minute he will finish 1/30 of the job. Do the same reasoning for what follows. Solve the above equation for x

OpenStudy (anonymous):

\[ \frac{1}{30}+\frac{1}{50}=\frac{1}{x}\\ 5 x + 3 x = 150\\ 8x = 150\\ x = \frac {75}4=18.75= 18 \text { min } 45 \text { sec } \]